OpenStudy (anonymous):

2 - 15n = 5 ( -3n + 2 )

hartnn (hartnn):

no solution

OpenStudy (theviper):

2 - 15n = 5 ( -3n + 2 ) 2-15n=-15n+10 \(2\neq10\) Therefore equation is wrong :)

OpenStudy (nubeer):

2-15n = -15n +10 2= 10 Solution not possible

OpenStudy (theviper):

Understood @Rubie #